600 is 5/9 of what number?
What percent of 50 is 40?
When Merideth cleaned her room she found the raio of clean clothes to dirty clothes was 2 to 5. If 35 articles of clothing were discovered how many were clean?
The diameter of a round skating rink is 14 m. Find the circumference of the rink to the nearest meter.
Simplify -7.2/6

Answer by stanbon(75887)

600 is 5/9 of what number?
(5/9)x = 600
x = 5400/5
x = 1080
----------------------------------------
What percent of 50 is 40?
x/100 = 40/50
x = 100(4/5)
x = 20*4
x = 80%
====================
When Merideth cleaned her room she found the raio of clean clothes to dirty clothes was 2 to 5. If 35 articles of clothing were discovered how many were clean?
-----
2x + 5x = 35
x = 5
clean clothes = 2x = 10 articles
=====================================
The diameter of a round skating rink is 14 m.
Find the circumference of the rink to the nearest meter.
circumference = (pi)d = 14pi = 43.982 meters
-----------------------------
Simplify -7.2/6 = -1.2
============
Cheers,
Stan H.
=============